# DONOD: Robust and Generalizable Instruction Fine-Tuning for LLMs via Model-Intrinsic Dataset Pruning

### v1

@@ Appendix A Appendix
A.1 Benchmark Details
@@
A.2 Training Settings
  • Global batch size: 16
  • Micro-batch size: 1
  • Learning rate: 2e-5
  • Optimizer: AdamW (Loshchilov & Hutter (2019))
  • Warmup-ratio 0.025
  • Every model is trained on 8 NVIDIA A100 GPUs for less than 1 hour
A.3 Datasets Details
@@
A.4 Cross Architecture Results
Table 6: Experimental results comparing different methods on Llama-2-13b-chat. The Non-Target column shows average performance excluding mathematical reasoning (target domain), revealing how methods generalize to other abilities. All values are percentages.
ALL (100%) 30.57 54.68 23.38 29.94 38.39 35.28
DONOD 20% 31.52 55.58 26.89 29.39 38.83 36.57
DONOD 30% 31.30 54.61 23.39 29.57 38.49 35.53
@@ A.4 Cross Architecture Results
A.5 Integration with TOPSIS
  1. 1.
    $\tilde{w}_{i,j}=\frac{w_{i,j}}{\sqrt{\sum_{k=1}^{n}w_{k,j}^{2}}}.$ (5)
  2. 2.
    $Z^{+}=[\max(\tilde{W}_{\text{DON}}),\min(\tilde{W}_{\text{NOD}})],\quad Z^{-}=[\min(\tilde{W}_{\text{DON}}),\max(\tilde{W}_{\text{NOD}})].$ (6)
  3. 3.
    $D_{i}^{+}=\sqrt{\sum_{j=1}^{2}(\tilde{w}_{i,j}-Z_{j}^{+})^{2}},\quad D_{i}^{-}=\sqrt{\sum_{j=1}^{2}(\tilde{w}_{i,j}-Z_{j}^{-})^{2}}.$ (7)
  4. 4.
    TOPSIS Score: The final score for each sample is:
    $C_{i}=\frac{D_{i}^{-}}{D_{i}^{+}+D_{i}^{-}},$ (8)
A.6 Ablation Configurations
@@
A.7 Bad Samples Gallery
How many edges does a cube have?
[…]
Find the number of vertices of the rectangular prism.
[…]
What is the smallest number among the three numbers 5, 8, and 4?
[…]
You subtracted 4.3 when you should have added 4.3 to a number, resulting in 8.8. Find the correct value.
[…]
Karen bakes 50 chocolate chip cookies. She keeps 10 for herself, and she gives 8 to her grandparents. Karen wants to give everyone in her class cookies. Each person in her class will receive 2 cookies. How many people are in Karens class?
Therefore the final answer is Karen has 16 people in her class.
Ronnie is a train engineer. His train pulls 3 red boxcars, 4 blue boxcars, and 7 black boxcars. The black boxcars can each hold 4000 pounds of coal, and the red boxcars can hold 3 times as much as the blue boxcars. All of the trains boxcars combined can hold 132,000 pounds of coal. What is the ratio of the coal capacity of each black boxcar to each blue boxcar?
I understand that Ronnies train has 3 red boxcars, 4 blue boxcars, and 7 black boxcars, and we need to find the ratio of the coal capacity of each black boxcar to each blue boxcar.I recognize that the black boxcars can hold 4000 pounds of coal, and the red boxcars can hold 3 times as much as the blue boxcars, and the total capacity of all boxcars is 132,000 pounds.I will let the coal capacity of each blue boxcar be B pounds, so each red boxcar can hold 3B pounds.I can write an equation based on the total capacity of all boxcars: 3(3B) + 4B + 7(4000) = 132,000.I simplify the equation by combining like terms: 9B + 4B + 28,000 = 132,000.I combine like terms to get: 13B + 28,000 = 132,000.I subtract 28,000 from both sides to get: 13B = 104,000, and then divide both sides by 13 to get: B = 8,000.I know that each blue boxcar can hold 8,000 pounds, and each black boxcar can hold 4,000 pounds, so the ratio of the coal capacity of each black boxcar to each blue boxcar is 4000/8000 = 1/2.
[…]
A.8 Empirical Validations of Output Layer Focus

### v2

@@ Appendix A Appendix
A.1 Theoretical Foundations
The generalization gap for a learning algorithm, for a specific function $f$, is the difference between its true (expected) risk $R_{\mathcal{D}}(f)$ over the data distribution $\mathcal{D}$ and its empirical risk $R_{S}(f)$ over a finite sample $S=\{(x_{1},y_{1}),\dots,(x_{m},y_{m})\}$. A smaller generalization gap means the model’s performance on unseen data is closer to its performance on training data, i.e., the model generalizes well.
The Rademacher complexity $R_{m}(\mathcal{F})$ of a hypothesis class $\mathcal{F}$ measures its ability to fit random noise. It is formally defined ((Bartlett and Mendelson 2003) and (Neyshabur, Tomioka, and Srebro 2015)) as:
$R_{m}(\mathcal{F})=\mathbb{E}_{\boldsymbol{\xi}\in\{\pm 1\}^{m}}\left[\frac{1}{m}\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{m}\xi_{i}f(x_{i})\right|\right]$
Similarly by (Bartlett and Mendelson 2003), for a hypothesis class $\mathcal{F}$ of real-valued functions and a 1-Lipschitz loss function $\ell$, a standard generalization bound, if inputs are bounded and output is bounded, states that for any $f\in\mathcal{F}$, with probability at least $1-\delta$ over the random draw of the sample $S$:
$R_{\mathcal{D}}(f)\leq R_{S}(f)+2R_{m}(\mathcal{F})+C\sqrt{\frac{\log(1/\delta)}{m}}$
where $C$ is a constant related to the range of the loss function, and for 1-Lipschitz losses, it simplifies to $2R_{m}(\mathcal{F})+\sqrt{\frac{\log(1/\delta)}{2m}}$. This inequality implies that the generalization gap $R_{\mathcal{D}}(f)-R_{S}(f)$ is bounded by terms that include the Rademacher complexity of the hypothesis class $\mathcal{F}$. Therefore, to show that the generalization gap of $M^{\prime}$ is smaller, we need to show that the Rademacher complexity of its corresponding hypothesis class is smaller.
A.2 Theoretical Analysis
Here, based on our case, we simplify our network as a network with a fixed architecture ($d$ layers, width $H$) and RELU activations, and consider a hypothesis class of functions whose weights $w$ satisfy a bound on $\mu_{p,q}(w)$. Specifically, we are interested in the Frobenius norm, which is $\mu_{2,2}(w)$. Let $\mathcal{F}_{\mu}=N_{d,H,\sigma_{RELU}}^{\mu_{2,2}\leq\mu}$ be the class of functions that can be realized by such a network where the overall $\ell_{2}$ norm of its weights is at most $\mu$.
For the model $M_{high}$ of weight $W_{high}$, its actual Frobenius norm is $\mu_{M_{high}}=\mu_{2,2}(W_{high})$. The function $f_{W_{high}}$ computed by $M_{high}$ belongs to the hypothesis class $\mathcal{F}_{\mu_{M_{high}}}$.
Similarly, the network $M_{low}$ has weights $W_{low}$, and its actual Frobenius norm is $\mu_{M_{low}}=\mu_{2,2}(W_{low})$. The function $f_{W_{low}}$ computed by $M_{low}$ belongs to the hypothesis class $\mathcal{F}_{\mu_{M_{low}}}$. We are given $\mu_{M_{low}}>\mu_{M_{high}}$.
Here, we theoretically prove that models with high DON generalize better.
To resonate the use of DON, we first introduce two key mathematical tools, the generalization gap and Rademacher complexity, as shown in Appendix 1.1. Let $M$, $M_{high}$ and $M_{low}$ be three neural networks with identical architectures but distinct weight matrices $W$, $W_{low}$ and $W_{high}$, respectively. Here $M_{high}$ stands for a model with high DON, $M_{low}$ stands for a model with low DON, and $M$ is an arbitrary auxiliary model.
We show that the model with higher DON generalizes better, e.g., if
$\begin{split}\text{DON}^{\text{high}}-\text{DON}^{\text{low}}&=\sum_{l=1}^{L}\left(\|W^{l}\|_{F}-\|W^{l}_{\text{high}}\|_{F}\right)\\ &\quad-\sum_{l=1}^{L}\left(\|W^{l}\|_{F}-\|W^{l}_{\text{low}}\|_{F}\right)>0\end{split}$ (5)
then $M_{high}$ generalize better.
Simplify 5, we have:
$\sum_{l=1}^{L}(\|W^{l}_{low}\|_{F}-\|W^{l}_{high}\|_{F})>0$ (6)
Using the same notation as (Neyshabur, Tomioka, and Srebro 2015), let $\mu_{2,2}(W_{low})=\sum_{l=1}^{L}\|W^{l}_{low}\|_{F}$ and $\mu_{2,2}(W_{high})=\sum_{l=1}^{L}\|W^{l}_{high}\|_{F}$, suppose the Frobenius norms of the weights satisfy $\mu_{2,2}(W_{low})>\mu_{2,2}(W_{high})$, we want to proof that that the generalization gap of $M_{high}$ is smaller than that of $M_{low}$.
Proof. Firstly, we make necessary assumptions and the setup based on our practical conditions as detailed in Appendix 1.2. According to the Corollary 2 of (Neyshabur, Tomioka, and Srebro 2015), for any $d\geq 1$, $1\leq p<\infty$, and $1\leq q\leq p^{*}=p/(p-1)$ (where $1/p+1/p^{*}=1$), the Rademacher complexity of the class $N_{\mu_{2,2}\leq\mu}^{d,H,\sigma_{RELU}}$ is bounded.
In our case, we are considering the Frobenius norm, so $p=2$ and $q=2$. This means $p^{*}=2/(2-1)=2$. Since $q=2$ and $p^{*}=2$, the condition $q\leq p^{*}$ ($2\leq 2$) is met. Therefore, we can apply the bound from Corollary 2:
$R_{m}(N_{\mu_{2,2}\leq\mu}^{d,H,\sigma_{RELU}})\leq\left(\frac{2\mu}{\sqrt[2]{d}}\right)^{d}R^{linear}_{m,2,D}$
Here, $R^{linear}_{m,2,D}$ is the Rademacher complexity of $D$-dimensional linear predictors with unit $\ell_{2}$ norm with respect to a set of $m$ samples. Since $p=2$, by the same Corollary 2 (Neyshabur, Tomioka, and Srebro 2015), we have a bound for this term: $R^{linear}_{m,2,D}\leq\sqrt{\frac{{\min{\{p^{*},4log(2D)}\}\max_{i}\|x_{i}\|^{2}_{p^{*}}}}{m}}$. Let $K=\frac{2^{d}}{d^{d/2}}\sqrt{\frac{{\min{\{p^{*},4log(2D)}\}\max_{i}\|x_{i}\|^{2}_{p^{*}}}}{m}}$. This $K$ is a positive constant that depends only on the fixed architecture ($d$), input dimensionality ($D$), sample size ($m$), and the maximum $\ell_{2}$ norm of input data points (which is assumed finite).
So, the Rademacher complexity bound for our class becomes:
$R_{m}(N_{d,H,\sigma_{RELU}}^{\mu_{2,2}\leq\mu})\leq K\cdot\mu^{d}$
For network $M_{high}$, its function $f_{W_{high}}$ belongs to the class $\mathcal{F}_{\mu_{M_{high}}}$, and its Rademacher complexity is bounded by:
$R_{m}(\mathcal{F}_{\mu_{M_{high}}})\leq K\cdot\mu_{M_{high}}^{d}$
For network $M_{low}$, its function $f_{W_{low}}$ belongs to the class $\mathcal{F}_{\mu_{M_{low}}}$, and its Rademacher complexity is bounded by:
$R_{m}(\mathcal{F}_{\mu_{M_{low}}})\leq K\cdot\mu_{M_{low}}^{d}$
As explained in Appendix 1.2, since we are given $\mu_{M_{low}}>\mu_{M_{high}}$, and $\mu_{M},\mu_{M^{\prime}}$ are non-negative (being norms), it directly follows that $\mu^{d}_{M_{low}}>\mu^{d}_{M_{high}}$. Therefore:
$R_{m}(\mathcal{F}_{\mu_{M_{high}}})\leq K\cdot\mu_{M_{high}}^{d}<K\cdot\mu_{M_{low}}^{d}$ (7)
This shows that the Rademacher complexity of the hypothesis class associated with $M_{high}$ is strictly smaller than that associated with $M_{low}$.
From the generalization bound established in Appendix A.1: For $M_{high}$, with probability at least $1-\delta$:
$R_{\mathcal{D}}(f_{W_{high}})-R_{S}(f_{W_{high}})\leq 2R_{m}(\mathcal{F}_{\mu_{M_{high}}})+C\sqrt{\frac{\log(1/\delta)}{m}}$
For network $M_{l}ow$, with probability at least $1-\delta$:
$R_{\mathcal{D}}(f_{W_{low}})-R_{S}(f_{W_{low}})\leq 2R_{m}(\mathcal{F}_{\mu_{M_{low}}})+C\sqrt{\frac{\log(1/\delta)}{m}}$
Since (7), it implies that the upper bound on the generalization gap for $M_{high}$ is lower than that for $M_{low}$, that is, in the worst case, the generalization gap $M_{high}$ can have is strictly less than $M_{low}$. In other words, $M_{high}$ is generally better generalized. We also provide a more intuitive interpretation of DON and NOD metrics in Appendix A.3.
A.3 Intuition Behind DON and NOD Metrics
DON as a Proxy for Generalization: A negative DON indicates that fine-tuning on $D_{i}$ increases the model’s weight magnitude. As shown above, we have proved that the model with high DON lead to better generalization. This is also supported by the study of (Shalev-Shwartz and Ben-David 2014), its increment relates to the regularization principles, where smaller norms often correlate with lower generalization error (e.g., weight decay). Intuitively, samples that induce a significant increase in the Frobenius norm contribute to the complexity of the model, potentially damage its ability to generalize across domains, and high and positive DON indicate the simplification of the model and better generalization. Our experimental results support this intuition, showing that samples yielding high DON positive values improve cross-domain accuracy.
NOD as an Indicator of Bad Sample: A high NOD value indicates that the data point $D_{i}$ has a significant influence on the model. In the context of ad-hoc SFT, training begins with a model that already possesses a certain level of generalization. Since the model is unlikely to encounter entirely new information or learn fundamentally new concepts after pretraining, data points with high NOD values are often indicative of low-quality samples, such as mislabeled or noisy data. Therefore, our method focuses on filtering out samples with high NOD values, thereby removing noisy or unlearnable data from the training process.
A.4 Choice of Frobenius Norm
A.5 Integration with TOPSIS
  1. 1.
    $\tilde{w}_{i,j}=\frac{w_{i,j}}{\sqrt{\sum_{k=1}^{n}w_{k,j}^{2}}}.$ (8)
  2. 2.
    $Z^{+}=[\max(\tilde{W}_{\text{DON}}),\min(\tilde{W}_{\text{NOD}})]$ (9)
    $Z^{-}=[\min(\tilde{W}_{\text{DON}}),\max(\tilde{W}_{\text{NOD}})]$ (10)
  3. 3.
    $D_{i}^{+}=\sqrt{\sum_{j=1}^{2}(\tilde{w}_{i,j}-Z_{j}^{+})^{2}}\quad$ (11)
    $D_{i}^{-}=\sqrt{\sum_{j=1}^{2}(\tilde{w}_{i,j}-Z_{j}^{-})^{2}}$ (12)
  4. 4.
    TOPSIS Score: The final score for each sample is:
    $C_{i}=\frac{D_{i}^{-}}{D_{i}^{+}+D_{i}^{-}},$ (13)
A.6 Training Settings
  • Global batch size: 16
  • Micro-batch size: 1
  • Learning rate: 2e-5
  • Optimizer: AdamW (Loshchilov and Hutter 2019)
  • Warmup-ratio 0.025
  • Every model is trained on 8 NVIDIA A100 GPUs for less than 1 hour
A.7 Bad Samples Gallery
1How many edges does a cube have?
[…]
3
4Find the number of vertices of the rectangular prism.
[…]
6
[…]
8Therefore the final answer is There are 22 multiples of 3 that can be formed by selecting three of the four numbers 1, 2, 3, and 4 (allowing duplicates) and arranging them in a row to form a three-digit number.
9
10What is the smallest number among the three numbers 5, 8, and 4?
[…]
12
[…]
15
16You subtracted 4.3 when you should have added 4.3 to a number, resulting in $8.$ Find the correct value.
[…]
18
19Karen bakes 50 chocolate chip cookies. She keeps 10 for herself, and she gives 8 to her grandparents. Karen wants to give everyone in her class cookies. Each person in her class will receive 2 cookies. How many people are in Karens class?
20Therefore the final answer is Karen has 16 people in her class.
21
22Ronnie is a train engineer. His train pulls 3 red boxcars, 4 blue boxcars, and 7 black boxcars. The black boxcars can each hold 4000 pounds of coal, and the red boxcars can hold 3 times as much as the blue boxcars. All of the trains boxcars combined can hold 132,000 pounds of coal. What is the ratio of the coal capacity of each black boxcar to each blue boxcar?
23I understand that Ronnies train has 3 red boxcars, 4 blue boxcars, and 7 black boxcars, and we need to find the ratio of the coal capacity of each black boxcar to each blue boxcar.I recognize that the black boxcars can hold 4000 pounds of coal, and the red boxcars can hold 3 times as much as the blue boxcars, and the total capacity of all boxcars is 132,000 pounds.I will let the coal capacity of each blue boxcar be B pounds, so each red boxcar can hold 3B pounds.I can write an equation based on the total capacity of all boxcars: 3(3B) + 4B + 7(4000) = 132,000.I simplify the equation by combining like terms: 9B + 4B + 28,000 = 132,000.I combine like terms to get: 13B + 28,000 = 132,000.I subtract 28,000 from both sides to get: 13B = 104,000, and then divide both sides by 13 to get: B = 8,000.I know that each blue boxcar can hold 8,000 pounds, and each black boxcar can hold 4,000 pounds, so the ratio of the coal capacity of each black boxcar to each blue boxcar is 4000/8000 = 1/2.
[…]
A.8 Empirical Validations of Output Layer Focus
Figure 11: Ranking of Frobenius norm delta of layers of Llama-2-13b-chat after fine-tuning on SAT Math COT dataset, epoch 2
A.9 Benchmark Details
@@
A.10 Datasets Details
@@
GSM8K Mathematical Problem-Solving COT 8794
Ifeval-Like Data Instruction Following General text 56.3K
A.11 Cross Architecture Results
@@
A.12 Ablation Configurations